OpenStudy (anonymous):

I am trying to convert 79/100 to an integer. Can someone help me

myininaya (myininaya):

you cant convert it to an integer

OpenStudy (anonymous):

if someone can help you do that, i would be impressed haha

myininaya (myininaya):

jf you gave a rational expression and the top is smaller than the bottom then it is definitely not an integer examples: 2/4=1/2 is not an integer; it is rational 4/2=2 is an integer and is also rational an integer is any number that can be written as one of the numbers in the set {...,-3,-2,-1,0,1,2,3,...}

myininaya (myininaya):

unless the top is 0 0/25 is an integer because 0/25=0

OpenStudy (anonymous):

Change the base to .1 and you get integer 79.